Use the gauss-Jordan method to solve the following system:
2x + y = 13
-2x + 3y = -1
How do I do this?

Answer by TimothyLamb(4379)

2x + y = 13
-2x + 3y = -1
---
put the system of linear equations into standard form
---
2x + y = 13
-2x + 3y = -1
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 5
y = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---